OpenStudy (anonymous):

Solve: x/4=8 a. 4 b. -4 c. 32 d. -32

OpenStudy (darkigloo):

multiply both sides by 4. x=32

OpenStudy (anonymous):

oh yeah i remember now thanks!

OpenStudy (squirrels):

Just multiply 4 to the other side dood

OpenStudy (anonymous):

wait would it be -32 though?

OpenStudy (squirrels):

Nupe

OpenStudy (whpalmer4):

\[\frac{x}{4} = 8\]Multiply both sides by 4\[4*\frac{x}{4} = 8*x\]\[\cancel{4}*\frac{x}{\cancel{4}} = 32\]\[x=32\]

OpenStudy (whpalmer4):

sorry, that should have been \[4*\frac{x}{4} = 8*4\]